OpenStudy (anonymous):

What is the sum of the solutions of 2{x-1}-4=-2 ? A. 1 1/2 B. 0 C. -2 D. 2

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

@Astrophysics

OpenStudy (zzr0ck3r):

is that \(2|x-1|-4=-2\) ?

OpenStudy (astrophysics):

Are those parentheses/brackets or absolute value as zz suggested?

OpenStudy (anonymous):

absolute like zz

OpenStudy (zzr0ck3r):

\(2|x-1|-4=-2\iff|x-1|-2=-1\iff|x-1|=1\iff x-1=1\text{ and } \\x-1=-1\)

OpenStudy (anonymous):

what?!?!?

OpenStudy (astrophysics):

\[2|x-1|-4=-2\] Well you could always plug in the solutions given and see if left side = right side, or you can solve for x. \[|x-1| = 1 \] \[x-1 = 1 \implies x = 2\]

OpenStudy (anonymous):

oh ok??

OpenStudy (astrophysics):

Alright I'll show you step by step,   technically your answer should also be x = 0, since we're dealing with absolute values, and I'm not entirely sure what you've been taught. But I hope this helps, good luck :)